A 446-g piece of copper tubing is heated to 89.5 d6Q11
Answer #1

Heat lost by tubing=Heat gained by water+vessel

446×0.387×(89.5-T)=159×1×(T-22.5)+10×(T-22.5)

15447.8-172.6T=169T-3802.5

T=56.35°C(approx).